OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7> -8.3° 1.7° 3.3° 13.3°

OpenStudy (phi):

do you know how to do a dot product ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8*9 + 7*7 72 + 49 119

OpenStudy (phi):

check your addition meanwhile, the other definition of dot product is \[ a \cdot b = |a| |b| cos\theta \] where |a| means length of a and | b | is length of b

OpenStudy (anonymous):

i mean 121, sorry wrong math

OpenStudy (phi):

so you now know \[ a \cdot b = 121 \] and \[ 121= |a| |b| \cos\theta \] you need to find |a| and |b|

OpenStudy (anonymous):

ok so sqrt(145) for a and sqrt(98) for b

OpenStudy (phi):

are you doing u = <8, 7>, v = <9, 7> ?

OpenStudy (phi):

|u|^2 is u dot u

OpenStudy (anonymous):

ok how do you find the length of a and of b, im not quite sure what i did but it must be wrong

OpenStudy (phi):

find |u|^2 = u dot u then take the square root do the same for v

OpenStudy (anonymous):

ok give me just a second to do that then

OpenStudy (phi):

let's use u and v and not a and b, to avoid confusion

OpenStudy (anonymous):

u^2 = 113 v^2 = 130

OpenStudy (phi):

yes, so |u| = sqrt(113) and |v|= sqrt(130)

OpenStudy (phi):

\[ 121= \sqrt{113} \sqrt{130} \cos\theta \] solve for theta

OpenStudy (anonymous):

.9983308057 = cosθ

OpenStudy (phi):

now take the inverse cosine of both sides

OpenStudy (phi):

they want the angle to the nearest 10th of a degree

OpenStudy (anonymous):

.0577868302

OpenStudy (phi):

I think you are in "radian mode" put your calculator in degree mode

OpenStudy (phi):

or multiply .0577868302 * 180/pi and round to the nearest tenth

OpenStudy (anonymous):

its about 3.3 degrees correct?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

and one more thing Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (anonymous):

for this one do i just multiply all the v and w vectors. then add?

OpenStudy (phi):

you multiply corresponding components and sum

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

is that write phi